Combustion analysis of a hydrocarbon produced 33.01g of CO2

and 5.41 g of H2O



1
Expert's answer
2021-10-05T01:45:57-0400

Moles of "CO_2=\frac{33.01}
{44.01 }=0.75"


Moles of "H_2O=\frac{5.41}
{18.01528}=0.3"


Raito "\frac{0.75}{0.3}=2.5" "\frac{0.3}{0.3}=1"



"C_2H_5"